OpenStudy (anonymous):

how can i change θ=pi/3 into rectangular form?

myininaya (myininaya):

Use the factor that arctan(y/x)=theta

OpenStudy (anonymous):

arctan (y/x)=pi/3 and then what?

myininaya (myininaya):

take tan of both sides what is tan(pi/3)?

OpenStudy (anonymous):

\[\sqrt{3}\]

myininaya (myininaya):

yes and so y/x= sqrt(3)

OpenStudy (anonymous):

thats it? thats the rectangular form?

myininaya (myininaya):

rectangular form is in terms of y and x if you want to multiply x on both sides you could